# Market Reaction to News Flows in Supply Chain Networks

### v3

@@ 4 Results and Discussion
In column (1) of Table 4, the estimated is 0.323 and highly significant, indicating that the probability of positive sentiment in a news article about a firm is positively correlated with the percentage change in the firm’s stock price from 2 days to 1 day before the disclosure of the news. Since the positiveness measure, in equation (3), ranges from 0 to 1 (Figure 3), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively () is 0.323% points higher than that of a firm mentioned with no positive sentiment () before the disclosure of the news. There are two possibilities for this positive coefficient. First, positive information included in news articles diffuses to stock markets partially through private channels before the formal disclosure of news, generating the pre-news effect. Second, stock prices of firms positively mentioned by news articles grow faster intrinsically than those of firms with no positive mention, leading to endogeneity biases.
@@
Column (1) of Table 4 further finds that the estimated is 0.923, indicating that the stock price of a firm mentioned in a news article most positively grows faster by 0.923% points than that of a firm mentioned with no positive sentiment after the news disclosure. The second and third rows from the bottom show and the value from a test for the null hypothesis that the difference is 0. In column (1), the value is close to 0, and thus we reject the null hypothesis. This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned with no positive sentiment even before the disclosure of news, positive news raises stock prices of former firms more by 0.6% points after its disclosure.
@@
These results in column (1) for the time window of 1 day generally hold in columns (2)-(5) for time windows of 2-5 and 30 days, while the post- and pre-news effects and the difference between the two decrease as the time window becomes wider. By contrast, the difference between the post- and pre-news effects is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size. This result implies that the intrinsic difference between firms mentioned positively and those with no positive mention in news articles is negligible.
@@
Panel (A) of Figure 5 graphically demonstrates the argument above. The left half of the figure where the value on the horizontal axis is negative shows the point estimate of the pre-news effect for different time windows, whereas the right half shows the post-news effect. Although the confidence intervals are added in the figure, they are too small to be recognized. The estimated coefficients show an inverted-U shape where the coefficients are close to 0 on the left and right edges. This shape suggests the followings. The change rate of the stock price of firms positively mentioned by news articles is not intrinsically different from that of firms negatively mentioned long before the disclosure of the news. However, the change rate of the stock price of positively mentioned firms began to rise compared to firms with no positive mention 30 days before the disclosure, peaking one day after the disclosure. Then, the difference in the change rates of stock prices between positively mentioned firms and those with no positive mention gradually decreases over time, converging to nearly zero 180 days after the disclosure.
The effect of the probability of negative sentiment in each news article shown in Panel (B) of Figure 5 is consistent with the effect of the probability of positive sentiment in Panel (A). The effect of negative sentiment in a news article is negligible 180 days or more before its disclosure but becomes negative and significant 30 days or fewer before the disclosure, implying information diffusion through private channels. After the disclosure, the negative effect becomes larger in the absolute term, suggesting a negative net effect of the disclosure of negative news on stock prices, in addition to the negative effect through private channels.
Indirect effect on suppliers and clients through global supply chains
@@
Next, we examine whether positive news about a firm affects stock prices of the firm’s suppliers and clients through global supply chains by estimating equation (3) where the positiveness index of news article about firm on day , , is replaced with the index for any supplier or client of firm .
Panels (A) and (B) of Figure 6 are the graphical presentation of the results for the effect of positive and negative news about firms, respectively, on their suppliers. Table 5 particularly shows the difference between the post- and pre-news effects of positive news and its values. The corresponding table for the effect of negative news is omitted here but is available in the Supplementary Information (SI Table 1). The results shown in the figure and table are similar to the results for the direct effect on own firms’ stock prices (Figure 5 and Table 4). The post- and pre-news effects of positive news are positive and significant, and their difference is positive and significant when the time window is narrow. These results suggest that positive news about firms raises stock prices of the firms’ suppliers before the disclosure of the news due to information diffusion through private channels and raises them more after the disclosure due to the additional effect of news medias. The effect of negative news articles is opposite in the direction and similar in size to the effect of positive ones, indicating the same conclusion for negative news.
@@
However, there are two notable differences between the direct effect on own firms and the indirect effect on suppliers. First, the post- and pre-news effect on suppliers and their difference is substantially smaller than the direct effects and their difference. For example, the pre-news effect on suppliers for the time window of 1 days is 0.0293 (Figure 6), while the corresponding direct effect is 0.323 (Figure 5), more than 10 times larger. Similarly, the net effect of the disclosure of news, represented by and shown in Table 5, is 0.033 when the time window is 1 days, while the corresponding value for the direct effect is 0.600 (Table 4). Therefore, we conclude that although positive news articles about firms affect stock prices of their suppliers, the effect on suppliers is less than 10% of the direct effect on own firms. Second, the difference between the post- and pre-news effects on suppliers is 0.033 when the time window is 1 day but becomes 0.004 when it is 5 days, meaning that the effect of the disclosure of positive news about a firm on its suppliers shrinks by 88% in 5 days. By contrast, its effect on the firm’s own stock price shrinks from 0.600 to 0.157 by 74%. This comparison implies that the net effect of the disclosure of news on suppliers lasts shorter after the disclosure than the effect on own stock prices.
Similarly, we examine the effect of positive and negative news articles on stock prices of client firms and illustrate the results in Panels (A) and (B) of Figure 7, respectively. Table 6 additionally presents the net effect of positive news, while the corresponding table for negative news can be found in the Supplementary Information (SI Table 2). These results are quite similar to those of the effect on suppliers shown in Figure 6 and Table 5. In summary, we find a positive (negative) and significant pre- and post-news effect of positive (negative) news and a positive (negative) and significant net effect of the disclosure of news. However, the effect of news about a firm on its clients is substantially smaller and lasts shorter than the effect on its own.
@@
We further use another sample that focuses on Japanese listed firms. The results for the direct effect of positive and negative news articles on Japanese firms’ own stock prices shown in Figure 8 are qualitatively and quantitatively similar to those using the sample of firms across the world shown in Figure 5. One subtle difference is that in Figure 8 for Japan, the post-news effect of negative news using the time window of 2 days is larger in the absolute term than that using the time window of 1 day, while Figure 5 for firms in the world shows the opposite. This finding suggests that the market reaction to negative news in Japan is slower than that to positive news in Japan and to positive and negative news in the world.
In addition, we examine the indirect effect of news articles about Japanese firms on stock prices of their suppliers and clients. A notable difference between this analysis on Japanese firms and the previous one on firms across the world is that more detailed domestic supply chains are identified based on firm-level surveys in the former than in the latter that mostly relies on financial statements (Data Section). Therefore, the average number of supply chain links for each Japanese firm in 2016 is 6.4, while that for each firm in the world is 3.3 (Tables 1 and 2).
@@
The results for the effect of positive and negative news on suppliers are demonstrated in Panels (A) and (B) of Figure 9, respectively, showing three notable differences from the previous results in Figure 6. First, the pre-news effect due to information diffusion through private channels is generally larger for Japanese firms than for others. For example, using the time window of 1 and 2 days, the pre-news effect of positive news on suppliers is 0.0090 and 0.0079 for Japan but 0.0029 and 0.0017 for others, respectively. Similarly, the post-news effect of negative news using the same windows is -0.0089 and -0.0066 for Japan and -0.0046 and -0.0025 for others, respectively.
@@
Second, the post-news effect of positive news is smaller than its pre-news effect (Panel [A] of Figure 9), although the post-news effect is always larger in size than the pre-news effect in the previous results. This finding implies that although positive information about a firm tends to diffuse actively through private channels, such as suggestions of financial companies, SNS, and word of mouth, before its disclosure by a news article and increase stock prices of suppliers of the firm. However, once the positive information is disclosed in a news article, the growth rate of the suppliers’ stock prices declines compared with that in the pre-news period. In other words, the net effect of the disclosure of positive news about a Japanese on stock prices of its suppliers is negative, although the previous results found the opposite.
Finally, we do observe that the post-news effect is larger than the pre-news effect in the case of negative news, as in the previous results (Panel [B] of Figure 9). However, the pre-news effect of negative news for the time window of 1 day is smaller in the absolute term than that for the longer time window of 2-5 days, although the previous result for firms across the world showed the opposite (Panel [B] of Figure 6). This, combined with the large negative post-news effect for the time window of 1 day, implies that the disclosure of negative information about a firm by a news article affects stock prices of its suppliers more substantially than diffusion of the information through private channels.
@@
We further estimate the effect on clients and show the results in Figure 10. Compared with the previous results for firms across the world shown in Figure 7, we find that the three notable differences found in the effect on suppliers still hold in the effect on clients: (1) The average effect is generally larger for Japanese firms than for firms across the world; (2) The net effect of the disclosure of positive news about a firm on stock prices of its clients is negative; (3) The pre-news effect of negative news for the time window of 1 day is smaller than that for longer time windows.
In summary, we find that the way positive and negative information about firms affects stock prices of their suppliers and clients before and after the disclosure of the information by news media is different between firms in Japan and those in other countries. Specifically, we find that positive news propagates more before its disclosure than after in the Japanese data. Although we cannot provide evidence for any mechanism behind the differences, one possible reason is the difference in the quality of supply chain information. The supply chain data used for firms worldwide in this study relies on financial statements, whereas the data for Japan is based on a questionnaire-based survey. Although the former is automatically determined by criteria based on sales volume, the latter identifies important trade partners, which are not determined only by sales volume. Since investors who can receive information about firms through private channels are more knowledgeable about supply chains and recognize the diffusion of performance through supply chains more accurately than other investors who rely on news media, once the knowledgeable investors receive positive information about a firm before its disclosure by news media, they invest in the firm’s suppliers and clients. This results in a large increase in stock prices of the suppliers and clients before the disclosure of the news. However, other investors who rely on news media and do not understand the role of supply chains react less strongly to positive news after its disclosure.
Another possible reason for this is the distinct nature of Japan’s supply chains, known as keiretsu. Links between suppliers and clients in these relationships are often stronger than those in other countries, as they are frequently associated with shareholding relationships, information exchanges, technical assistance, and research collaboration [67, 68]. These strong relationships may explain why positive news about firms may have a greater impact on their suppliers and clients. Because investors recognize the strength of these ties and anticipate the transmission of performance through supply chains, news about firms tends to have a greater impact on the stock prices of their suppliers and clients in Japan compared to other countries. However, these two reasons cannot explain why negative news does not propagate before disclosure, which should be investigated in future research.
@@ 5 Conclusion
This study examines how positive and negative news about firms affects the change rate of their stock prices and, moreover, how it affects the change rate of stock prices of the firms’ suppliers and clients, using a large sample of publicly listed firms across the world and another of Japanese listed firms. The level of positiveness and negativeness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements and supplemented by other information on the website and news medias, while those of Japanese firms are taken from large-scale firm-level surveys.
We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through private channels, such as SNS and word of mouth. Positive news also raises stock prices of the firms’ suppliers and clients before and after its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and clients than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for information diffusion through private channels. However, the post-news effect on suppliers and clients in Japan can be smaller than the pre-news effect, a result opposite to those from firms across the world. This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
@@
The data that support the findings of this study are available from FactSet Research Systems Inc. and Tokyo Shoko Research, Ltd, and London Stock Exchange Group Data & Analytics. but restrictions apply to the availability of these data, which were used under license for the current study, and so are not publicly available. Data are however available from the authors upon reasonable request and with permission of FactSet Research Systems Inc. and Tokyo Shoko Research, Ltd, and London Stock Exchange Group Data & Analytics.

### v4

@@ 4 Results
In column (1) of Table 4, the estimated is 0.323 and highly significant, indicating that a higher value of the positive sentiment index in a news article about a firm is associated with the percentage change in the firm’s stock price from 2 days to 1 day before the release of the news. Since the positiveness measure, in equation (5), ranges from 0 to 1 (Figure 3), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively () is 0.323% points higher than that of a firm mentioned with no positive sentiment () before the disclosure of the news. There are two possibilities for this positive coefficient. First, positive information included in news articles may be incorporated into stock prices prior to formal disclosure through non-public channels, contributing to pre-news price movements. Second, stock prices of firms positively mentioned by news articles may exhibit higher intrinsic growth than those of firms with no positive mention, leading to endogeneity biases. In the following analysis, we do not attempt to disentangle these two channels. Instead, we focus on the difference between post- and pre-news coefficient estimates, which summarizes the additional change in stock prices around public news disclosure relative to pre-existing price movements.
@@
Column (1) of Table 4 further shows that the estimated is 0.923, indicating that, in the post-news period, firms mentioned with positive news sentiment experience stock price growth that is higher by 0.923 percentage points than firms mentioned with no positive sentiment. The second and third rows from the bottom show and the value from a test for the null hypothesis that the difference is 0. In column (1), the value is close to 0, indicating that the difference is statistically different from zero. This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned with no positive sentiment even before the disclosure of news, stock prices of firms with positive news sentiment are higher by an additional 0.6% after its disclosure.
@@
These results in column (1) for the time window of 1 day generally hold in columns (2)-(6) for time windows of 2-5 and 30 days, while the post- and pre-news coefficient estimates and the difference between the two decrease as the time window becomes wider. By contrast, the difference between the post- and pre-news coefficient estimates is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size. This result implies that the intrinsic difference between firms mentioned positively and those with no positive mention in news articles is negligible.
@@
Panel (A) of Figure 5 graphically demonstrates the argument above. The left half of the figure where the value on the horizontal axis is negative shows the point estimate of the pre-news coefficient for different time windows, whereas the right half shows the post-news coefficient. Although the confidence intervals are added in the figure, they are too small to be recognized. The estimated coefficients show an inverted-U shape where the coefficients are close to 0 on the left and right edges. This pattern indicates the following. The change rate of the stock price of firms mentioned with positive news sentiment is not intrinsically different from that of firms mentioned with no positive news sentiment long before the disclosure of the news. However, the change rate of the stock price of positively mentioned firms began to rise compared to firms with no positive mention 30 days before the disclosure, peaking one day after the disclosure. Then, the difference in the change rates of stock prices between positively mentioned firms and those with no positive mention gradually decreases over time, converging to nearly zero 180 days after the disclosure.
The pattern of coefficients for negative sentiment shown in Panel (B) of Figure 5 is qualitatively consistent with the pattern of positive sentiment shown in Panel (A). The corresponding detailed regression results are reported in the Supplementary Information (SI Table 3). The coefficients are negligible 180 days or more before its disclosure but become negative and significant 30 days or fewer before the disclosure. After disclosure, the magnitude of the coefficients increase further.
Associations for suppliers and clients along global supply chains
@@
Next, we examine whether news sentiment about a firm is associated with stock prices of the firm’s suppliers and clients along global supply chains by estimating equation (5) where the positiveness or negativeness index of news article about firm on day , , is replaced with the index for any supplier or client of firm .
Panels (A) and (B) of Figure 6 present coefficient estimates for positive and negative news sentiment about firms and the associated stock price changes of their suppliers. Table 5 reports the difference between post- and pre-news coefficient estimates for positive sentiment, with corresponding results for negative sentiment reported in SI Table 1. Detailed regression results are reported in SI Tables 4 and 5.
For suppliers, both pre- and post-news coefficient estimates are positive and statistically significant for positive news sentiment, with a larger post–pre difference for shorter time windows. The corresponding coefficient estimates for negative sentiment exhibit the opposite sign and similar magnitudes. Overall, the magnitude of supplier-level coefficient estimates is substantially smaller than that observed for own-firm stock prices.
@@
However, there are two notable differences between the own-firm coefficients and those for suppliers. First, the post- and pre-news coefficients for suppliers and their difference are substantially smaller than the corresponding own-firm coefficients and their difference. For example, the pre-news coefficient for suppliers for the time window of 1 day is 0.0293 (SI Table 4), while the corresponding coefficient for own firms is 0.323 (Table 4), more than 10 times larger. Similarly, the difference between the post- and pre-news coefficients, , is 0.033 for suppliers when the time window is one day (Table 5), while the corresponding difference for own firms is 0.600 (Table 4). Quantitatively, although positive news articles about firms are associated with stock prices of their suppliers, the corresponding coefficients for suppliers are less than 10% of those for own firms. Second, the difference between the post- and pre-news coefficients on suppliers is 0.033 when the time window is 1 day but becomes 0.004 when it is 5 days, indicating that the post-pre difference in coefficients for suppliers declines by 88% in 5 days. By contrast, the coefficients for own firm decline from 0.600 to 0.157 by 74%. This comparison shows that the difference between the post- and pre-news coefficients for suppliers diminishes more quickly after disclosure than the corresponding difference for firms’ own stock prices.
As in the analysis for firms’ own stock prices, negative sentiment shows a pattern similar to positive sentiment for suppliers (SI Table 5). Both the post- and pre-news coefficients for suppliers, and their difference, are substantially smaller than the corresponding coefficients for firms’ own stock prices, and the post–pre difference for suppliers diminishes more quickly after disclosure.
Similarly, we examine coefficient estimates for positive and negative news sentiment in relation to stock prices of client firms and present the results in Panels (A) and (B) of Figure 7, respectively. We also show the difference between the post- and pre-news coefficient estimates for positive (Table 6) and negative (SI Table 2). Detailed regression results for clients are reported in SI Table 6 for positive sentiment and in SI Table 7 for negative sentiment. These results exhibit the same qualitative pattern as those for suppliers shown in Figure 6 and Table 5. Specifically, we find positive (negative) and significant pre- and post-news coefficients for positive (negative) news sentiment, as well as a positive (negative) and significant post–pre difference. However, the coefficients for clients are substantially smaller and decline more rapidly than those for own firms.
@@
We further use another sample that focuses on Japanese listed firms. The results for the own-firm coefficients of positive and negative news sentiment for Japanese firms’ own stock prices shown in Figure 8 are qualitatively and quantitatively similar to those using the sample of firms across the world shown in Figure 5. The detailed regression results are shown in SI Tables 8 and 9. One subtle difference is that in Figure 8 for Japan, the post-news coefficient for negative news sentiment using the time window of 2 days is larger in absolute terms than that using the time window of 1 day, while Figure 5 for firms in the world shows the opposite. This finding suggests that the market reaction to negative news in Japan is slower than that to positive news in Japan and to positive and negative news in the world.
In addition, we examine coefficients for news sentiment about Japanese firms on stock prices of their suppliers and clients. Detailed regression results for suppliers are reported in SI Tables 10 and 11, and those for clients are reported in SI Tables 12 and 13.
A notable difference between the Japanese supply chain data and the global supply chain data is that the former provides dense coverage of domestic supply chain links based on firm-level surveys, whereas the latter identifies major customer–supplier links primarily from financial statements (see Data Section). Consistent with this difference in coverage, the average number of supply chain links per listed firm in 2016 is 6.4 in Japan and 3.3 in the global sample (Tables 1 and 2).
@@
The coefficient estimates for positive and negative news sentiment in relation to suppliers’ stock prices are shown in Panels (A) and (B) of Figure 9, respectively, revealing three notable differences from the corresponding global results shown in Figure 6. First, the pre-news coefficients, which are larger in magnitude prior to disclosure, are generally larger for Japanese firms than for firms in the global sample. For example, for the one- and two-day time windows, the pre-news coefficients for positive news on suppliers are 0.0090 and 0.0079 in Japan, compared with 0.0029 and 0.0017 in the global sample. Similarly, for the same time windows, the post-news coefficients for negative news sentiment are -0.0089 and -0.0066 in Japan, compared with -0.0046 and -0.0025 in the global sample.
@@
Second, the post-news coefficient for positive news sentiment is smaller than its pre-news coefficient (Panel (A) of Figure 9), although the post-news coefficient is always larger in size than the pre-news coefficient in the global sample. This finding implies that positive information about a firm may be incorporated into prices prior to disclosure, for example suggestions of financial companies, SNS, and word of mouth, and that such information is associated with higher stock prices of the firm’s suppliers. However, once the positive information is disclosed in a news article, the magnitude of the coefficients for suppliers is smaller than that in the pre-news period. In other words, for Japanese firms, the difference between the post- and pre-news effects on suppliers is negative, indicating that the pre-news effect exceeds the post-news effect.
In contrast to the results for positive news, the response to negative news in Japan exhibits the same qualitative pattern as that observed in the global sample. Specifically, for negative news, the post-news effect is larger than the pre-news effect for Japanese firms, as is also the case for firms worldwide (Panel (B) of Figure 9 and Panel (B) of Figure 6).
Within this common pattern, however, there are differences across time windows. In particular, for Japanese firms, the pre-news effect of negative news for the one-day time window is smaller in absolute value than that for longer windows of two to five days, whereas the global results show an increase in the absolute value of the pre-news effect as the window size decreases.
@@
We further examine coefficients for news sentiment on client firms and present the results in Figure 10. Detailed regression results are reported in SI Tables 12 and 13. Compared with the corresponding results for firms across the world shown in Figure 7, we find that the same three differences observed for suppliers also apply to clients. First, the average magnitude of the effects is generally larger for Japanese firms than for firms in the global sample. Second, for positive news, the difference between the post- and pre-news effects on clients is negative in Japan. Third, for negative news, the pre-news effect for the one-day time window is smaller than that for longer time windows.
5 Discussion
In summary, we find that the way positive and negative information about firms affects stock prices of their suppliers and clients before and after the disclosure of the information by news media is different between firms in Japan and those in other countries. Additionally, we find that stronger pre-news stock price responses are observed for positive news in the Japanese sample.
It is important to emphasize that these results do not imply that investors possess detailed knowledge of firm-level supply chain structures. Rather, the observed stock price responses are consistent with market pricing that reflects information correlated with supply chain relationships. Several factors may be correlated with the observed stock price responses along supply chain relationships. The present analysis does not identify the mechanisms underlying these correlations.
It is also important to distinguish between realized economic impacts and market-based expectations reflected in stock prices. The present study focuses on stock price responses to news, which capture investors’ expectations about potential spillovers along supply chain relationships rather than confirmed changes in production, orders, or real economic outcomes. Whether such expectations are eventually realized in the form of actual production disruptions or demand reductions is beyond the scope of this analysis.
In this context, we clarify the specific dimension of market responses examined in this study. The focus of this study is on stock price responses, which reflect changes in market valuation. While other market indicators such as trading volume or volatility capture important aspects of market activity and uncertainty, they represent different dimensions of market behavior and are not direct measures of valuation effects. Accordingly, the present analysis concentrates on price-based responses to quantify how news-related information propagates through supply chain networks. Analyses of news spillovers in trading volume or volatility would require a different empirical design and are therefore not pursued in this study.
With this focus on price-based responses in mind, we turn to the interpretation of differences between the global and Japanese samples. These differences between the global and Japanese samples do not imply that different databases mechanically generate different results. Rather, the results consistently show that stock price responses propagate through supply chain relationships in both samples, while the timing and relative magnitude of such propagation differ across institutional contexts.
The use of different supply chain databases reflects these institutional contexts rather than causing the observed differences. The FactSet database captures major supply chain links disclosed in financial statements, whereas the TSR database identifies economically important trading partners through firm-level surveys. These databases therefore provide complementary views of supply chain relationships, rather than alternative measurements of the same underlying network.
Although the FactSet database is mainly determined by criteria based on sales volume, the TSR database identifies important trade partners, which are not determined only by sales volume. One possible interpretation is that the observed differences are consistent with heterogeneous timing and magnitude of market responses across institutional contexts. The present analysis does not establish the mechanisms generating these patterns.
Several institutional and historical factors may potentially contribute to the observed differences between Japan and other countries. The observed differences arise in an institutional context in which supply chain relationships in Japan are often discussed using the term keiretsu [3, 50]. The present analysis does not identify whether, or how, such institutional features contribute to the observed patterns. Other factors, such as historical shocks to domestic supply chains or differences between domestic and global supply chain structures, may also play a role. Identifying the relative importance of these mechanisms is beyond the scope of this study and remains an important direction for future research.
@@ 6 Conclusion
We find that news sentiment, both positive and negative, is associated with systematically higher and lower change rates of stock prices of firms mentioned in the news in the pre-news period, which may be consistent with a combination of anticipatory market behavior, information correlated with forthcoming news, and selection in news coverage. Such news sentiment is also associated with stock price responses of the firms’ suppliers and clients in both the pre-news and post-news periods, consistent with systematic valuation comovements along supply-chain relationships. In addition, we generally find that post-news effects on stock prices of the mentioned firms and their suppliers and clients are larger than pre-news effects. The difference between post- and pre-news effects summarizes differences in stock price changes between the post- and pre-disclosure periods. However, for Japanese firms, post-news effects on suppliers and clients can be smaller than pre-news effects, in contrast to the pattern observed for firms across the world.
@@
The data used in this study are obtained from commercial databases, including FactSet Research Systems Inc., Tokyo Shoko Research, Ltd., and London Stock Exchange Group Data & Analytics (Refinitiv). These data are subject to licensing restrictions and cannot be made publicly available by the authors. Researchers with access to the same commercial data sources under appropriate licenses can obtain the data directly from the respective providers and reproduce the analyses conducted in this study. The authors do not have special access privileges to these data beyond standard institutional licenses.